OpenStudy (safari321):

what is -square root 1000 + 8 square root 10

OpenStudy (anonymous):

\[-\sqrt{1000}+8\sqrt{10}\]\[=-\sqrt{100*10}+8\sqrt{10}\]\[=-10\sqrt{10}+8\sqrt{10}\]\[=(8-10)\sqrt{10}\]\[=-2\sqrt{10}\]

OpenStudy (safari321):

thanks